OpenStudy (jkristia):

find limit of sin(x) / .... I'm stuck trying to find the limit of (sin(x) - 1/2) / (x - pi/6) for x -> pi/6 algebraically. Any hints ? (please do not give the solution, just a hint of how to solve it)

OpenStudy (jkristia):

A bit easier to read \[\lim_{x \rightarrow \frac{\pi}{6}} \frac{ \sin(x) - \frac{1}{2}}{x - \frac{\pi}{6}}\]

OpenStudy (raden):

because it would be 0/0, u can use differentiate

OpenStudy (jkristia):

I have to find the limit without using L'Hospital. I know the the answer is \[\frac{\sqrt{3}}{2}\], but I can't figure out how to find it algebraically

OpenStudy (raden):

well, i will figure out for u